# Law of Large Numbers for continuous $N$-particle ensembles at fixed temperature

### v1

@@ 4 Sufficient conditions for the LLN
In this section, we assume that the sequence $\{\mu_{N}\in\mathcal{M}_{N}\}_{N\geq 1}$ is $\theta$-LLN-appropriate, in the sense that, if we let
@@
We shall prove that the numbers $m_{1},m_{2},\dots\in\mathbb{R}$, defined by the equation (13), satisfy the conditions of Definition 3.1. By virtue of Proposition 2.3, this final goal of the section is equivalent to proving
@@ 4.1 Technical lemmas on the action of Dunkl operators
Define the $k$-truncation polynomial
@@
proving the desired Eqn. (20).
If $\ell(\lambda)=2$, then $\lambda=(n_{1},n_{2})$, for some positive integers $n_{1}\geq n_{2}$. As above, we deduce that $\mathcal{D}_{i}^{n_{2}}\mathcal{D}_{j}^{n_{1}}e^{F_{N,\theta}(\vec{x})}\big|_{\vec{x}=(0^{N})}=\mathcal{D}_{2}^{n_{2}}\mathcal{D}_{1}^{n_{1}}e^{F_{N,\theta}(\vec{x})}\big|_{\vec{x}=(0^{N})}$, whenever $i\neq j$. Together also with (21), we obtain
@@
$N^{-|\lambda|-\ell(\lambda)}\left[\prod_{i=1}^{\ell(\lambda)}{\mathcal{P}_{\lambda_{i}}}\right]e^{F_{N,\theta}(\vec{x})}\bigg|_{\vec{x}=(0^{N})}=\left(1+O\Big(\frac{1}{N}\Big)\right)\cdot N^{-|\lambda|}\mathcal{D}_{1}^{\lambda_{1}}\cdots\mathcal{D}_{\ell(\lambda)}^{\lambda_{\ell(\lambda)}}e^{F_{N,\theta}(\vec{x})}\big|_{\vec{x}=(0^{N})}\\ +\sum_{j=1}^{\ell(\lambda)-1}\sum_{\begin{subarray}{c}k_{1},\dots,k_{j}\in\mathbb{Z}_{\geq 1}\\ k_{1}+\dots+k_{j}=|\lambda|\end{subarray}}O\Big(\frac{1}{N}\Big)\cdot N^{-k_{1}-\,\dots\,-k_{j}}\mathcal{D}_{1}^{k_{1}}\cdots\mathcal{D}_{j}^{k_{j}}e^{F_{N,\theta}(\vec{x})}\big|_{\vec{x}=(0^{N})}$
@@ 4.2 Proof of the “if” part of the main theorem
By Lemma 4.2 and the commutativity of Dunkl operators, the desired Equation (17) will follow if, for any $s\in\mathbb{Z}_{\geq 1}$ and arbitrary $\mu_{1},\dots,\mu_{s}\in\mathbb{Z}_{\geq 1}$, we manage to prove
$\lim_{N\to\infty}{N^{-|\mu|}\Big[\mathcal{D}_{s}^{\mu_{s}}\cdots\mathcal{D}_{2}^{\mu_{2}}\mathcal{D}_{1}^{\mu_{1}}\Big]e^{F_{N,\theta}(\vec{x})}\,\Big|_{\vec{x}=(0^{N})}}=\prod_{i=1}^{s}{m_{\mu_{i}}}.$ (23)

### v2

@@ 4 Sufficient conditions for the LLN
In this section, we assume that the sequence $\{\mu_{N}\in\mathcal{M}_{N}^{>0}\}_{N\geq 1}$ is $\theta$-LLN-appropriate, in the sense that, if we let
@@
The goal of this section is to prove that the numbers $m_{1},m_{2},\dots\in\mathbb{R}$, defined by equation (13), satisfy the conditions of Definition 3.1. By virtue of Proposition 2.3, this is equivalent to proving
@@ 4.1 Technical lemmas on the action of Dunkl operators
Define the $k$-truncated polynomial
@@
proving that the desired Eqn. (20) follows at once from Eqn. (19) for $\mu=(n)$.
If $\ell(\lambda)=2$, then $\lambda=(n_{1},n_{2})$, for some positive integers $n_{1}\geq n_{2}$. As above, we deduce that $\mathcal{D}_{i}^{n_{2}}\mathcal{D}_{j}^{n_{1}}e^{F_{N,\theta}(\vec{x})}\big|_{\vec{x}=(0^{N})}=\mathcal{D}_{2}^{n_{2}}\mathcal{D}_{1}^{n_{1}}e^{F_{N,\theta}(\vec{x})}\big|_{\vec{x}=(0^{N})}$, whenever $i\neq j$. Together with (21), we obtain
@@
$N^{-|\lambda|-\ell(\lambda)}\left[\prod_{i=1}^{\ell(\lambda)}{\mathcal{P}_{\lambda_{i}}}\right]e^{F_{N,\theta}(\vec{x})}\bigg|_{\vec{x}=(0^{N})}=\left(1+O\Big(\frac{1}{N}\Big)\right)\cdot N^{-|\lambda|}\mathcal{D}_{1}^{\lambda_{1}}\cdots\mathcal{D}_{\ell(\lambda)}^{\lambda_{\ell(\lambda)}}e^{F_{N,\theta}(\vec{x})}\big|_{\vec{x}=(0^{N})}\\ +\sum_{j=1}^{\ell(\lambda)-1}\sum_{\begin{subarray}{c}k_{1},\dots,k_{j}\in\mathbb{Z}_{\geq 1}\\ k_{1}+\dots+k_{j}=|\lambda|\end{subarray}}O\Big(\frac{1}{N}\Big)\cdot N^{-k_{1}-\,\dots\,-k_{j}}\mathcal{D}_{1}^{k_{1}}\cdots\mathcal{D}_{j}^{k_{j}}e^{F_{N,\theta}(\vec{x})}\big|_{\vec{x}=(0^{N})},$
@@ 4.2 Proof of the “if” part of the main theorem
By Lemma 4.2 and the commutativity of Dunkl operators, the desired equation (17) will follow if, for any $s\in\mathbb{Z}_{\geq 1}$ and $\mu_{1},\dots,\mu_{s}\in\mathbb{Z}_{\geq 1}$, we manage to prove
$\lim_{N\to\infty}{N^{-|\mu|}\Big[\mathcal{D}_{s}^{\mu_{s}}\cdots\mathcal{D}_{1}^{\mu_{1}}\Big]e^{F_{N,\theta}(\vec{x})}\,\Big|_{\vec{x}=(0^{N})}}=\prod_{i=1}^{s}{m_{\mu_{i}}}.$ (23)